OpenStudy (anonymous):

Quick quesiton: when multiplying -4x*8x^3 do I add the x's or would that be -32x^3?

OpenStudy (luffingsails):

x^n * x^m = x^(m+n) A simple example 3^2 * 3^3 = (3*3) * (3*3*3) = (3*3*3*3*3) = 3^(2+3) = 3^5 So, in your example: -4x*8x^3 = -32*x^4

OpenStudy (anonymous):

ah okay That's what I thought